# Personalize to generalize: Towards a universal medical multi-modality generalization through personalization

### v1

@@ 3.2.1 The connection between the constraints and global biological prior
It is important to note that the above constraints are closely interconnected, as they align with Eq. 1. After obtaining additional knowledge from $\mathbb{O}$, the invariance constraint ensures that the representations from each modality for a given individual are the same, such that ${\mathbb{X}^{i}_{h}}^{\prime}$ and ${\mathbb{X}_{h}}^{\prime}$ can be considered equivalent. Combined with the decomposition constraint, which enforces that ${\mathbb{X}_{h}}^{\prime}$ is shared for the generation of all possible modalities, ${\mathbb{X}_{h}}^{\prime}$ is thus able to generalize across modalities.
Additionally, the equivariance and decomposition constraints implicitly maintain SO(3)-equivariance by satisfying the relation $\mathcal{D}\circ\mathcal{E}(\phi^{i}(X^{i}_{h});\theta)=\phi^{i}(\mathcal{D}\circ\mathcal{E}(X^{i}_{h};\theta))$ [^1] The SO(3) transformations are left-multiplication; they are expressed here in a simplified form, using $\phi^{i}(\cdot)$., where $\theta$ represents the model parameters after training. This ensures that geometric transformations are preserved in the latent features $z^{i}_{h}$, such that $\mathcal{F}(z^{i}_{h})=\phi^{i}$. The invariance constraint then requires that $z^{i}_{h}$ with geometric transformations can retrieve features from $\mathbb{O}$ to form ${\mathbb{X}^{i}_{h}}^{\prime}$, which remains invariant to any geometric transformation. This implicitly constrains $\mathbb{O}$ to contain comprehensive biological information, including other potential geometric transformations, thereby improving the $\mathbb{X}_{h}^{\prime}$ through $\mathbb{O}$ in approximating $\mathbb{X}_{h}$ and enhancing the robustness of ${\mathbb{X}_{h}}^{\prime}$.
3.3 Application of different modalities and tasks
@@
Empirically, we adopt the SwinUNETR architecture [17] as the backbone of the encoder $\mathcal{E}$, and implement the proposed components. The model is trained with $\mathcal{L}_{pre}$ during the pre-training phase; users have the option to either use the standard SwinUNETR by loading only our pre-trained encoder weights or to employ our proposed model structure with all pre-trained weights for downstream tasks. Notably, all modalities for a given individual, $X_{h}={X^{i}_{h},X^{j}_{h},...,X^{k}_{h}},i,j,...,k\in\mathcal{M}$, share the same encoder, with the encoder’s channel size set to match the number of modality types. The input volume size for all experiments is fixed at $96\times 96\times 96$. Further empirical details on how $\mathbb{X}_{h}$ is leveraged for homogeneous and heterogeneous generalization are provided in Secs. 4 and 5.
4 Homogeneous generalization: MRI
This section demonstrates that our approach enhances the homogeneous generalization across structural modalities in MRI. To validate that our method captures personalized information, especially anatomical structure features, during the pre-training stage, we first apply it to modality transfer tasks. Next, we adapt the pre-trained model for the downstream missing modality segmentation task. Experimental results indicate that our approach outperforms state-of-the-art (SOTA) methods in both tasks, thereby supporting the $\mathbb{X}_{h}$ Hypothesis and confirming the effectiveness of our method.
4.1 Pre-training for modality transfer
Modality transfer tasks focus on converting medical images from multiple modalities to other modalities. We test our approach on the structural modalities of MRI. This task aligns seamlessly with our pre-training objective, and the quality of the generated modalities serves as a validation of the anatomical knowledge captured by our pre-trained model. Importantly, our approach can generate all modalities without knowing the exact modalities where the input form, as the learned representation ${\mathbb{X}_{h}}^{\prime}$ encompasses comprehensive information across all possible modalities.
@@ 4.1 Pre-training for modality transfer
Ours 87.63 3.25 91.19 89.49 1.82 87.45 1.03 85.17 3.93
Table 3: Missing modality segmentation results of MRI on BRATS18: Num denotes the number of missing modalities for different settings. We report mean and standard deviations of DICE results of all experimental combinations under the same Num. The best results are highlighted in blue. Please refer to Appendix Tab. 10 for detailed results of each setting. MN: Missing Num; FM: Full Modality.
@@
Experimental settings. Following previous methods [25], we utilize the multi-modal brain tumor segmentation challenge 2023 (BRATS23) dataset [1, 32, 3, 2]. BRATS23 includes four structural MRI modalities (T1, T1ce, T2, and FLAIR) for each individual. Our model is tested on the BRATS23 validation set, which contains these four modalities for 219 individuals. We evaluate the quality of synthesis using peak signal-to-noise ratio (PSNR), normalized mean squared error (NMSE), and structural similarity index (SSIM) [48]. To provide comprehensive results, we separately compare the translation results for T1 $\to$ T2 and T2 $\to$ FLAIR, as some previous methods are only capable of single-modality transfer. These include both 2D and 3D generation methods, as shown in Tab. 1. Additionally, we employed SwinUNTER for multi-modality translation comparisons. All evaluations were performed on 3D volumes; for the 2D methods, synthesized target images were stacked to form a 3D volume for comparison. Please refer to model training details to Appendix D.
@@
Results. Tab. 1 exhibits the transfer results of T1 $\to$ T2 and T2 $\to$ Flair. Our approach significantly surpasses previous 2D and 3D generation methods, including single- and multi-modality translation methods. Specifically, our approach exceeds current SOTA diffusion-based methods, such as 2D-based MD-Diff and 3D-based MS-SPADE. In terms of multi-modality translation, as Tab. 2 shows, our approach performs better than MS-SPADE and SwinUNETR across all metrics under all settings. Moreover, for all settings, it can be seen that our method significantly improves the SSIM, indicating a better anatomy structure obtained by our approach. These results indicate that the $\mathbb{X}_{h}$ Hypothesis is plausible for homogeneous generalization, and our personalized approach is able to obtain its approximation.
Method Dice↑ Dice-↑ TPR↑ TNR↑ FNR↓ FPR↓
From scratch
nnUnet [21] 33.10 - - - - -
UNETR [18] 10.81 23.14 80.65 3.64 19.35 96.36
SwinUNETR [17] 43.45 62.60 90.32 62.73 9.68 37.27
SwinUNETR with pre-train
With pre-train [40] 44.06 57.79 89.25 73.64 10.75 26.36
With ours 48.20 61.16 88.17 77.27 11.83 22.72
4.2 Tuning for missing modality segmentation
@@
Experimental settings. To validate the generalization ability of the pre-trained model, we fine-tune the model obtained from Sec. 4.1 on the BRATS18 [32] from the Multimodal Brain Tumor Segmentation Challenge. Similar to BRATS23, BRATS18 also consists of the same four structural modalities. We employ the Dice similarity coefficient (DICE) as the metric for evaluation. For a fair comparison, we follow data splits of [38] and reproduce the results of previous methods  on these splits by using their released code and following their original settings  [^2] Though we tried our best, it can be noticed some reproduced results are lowered than their reported results in their original paper. It should be clarified that our results also exceed those reported results. However, for a comprehensive study, we mainly report our reproduced results. . Additional experimental details can be seen in Appendix D.
@@
Results. Tab. 3 presents the segmentation results of our approach compared to previous methods. We also compute the standard deviation of DICE scores under various missing modality settings, which highlights the robustness of our model. Notably, our approach outperforms previous methods in most missing modality scenarios, particularly when the number of missing modalities is large. Our method shows significant improvements in enhancing tumor segmentation, especially when Missing Num = 3. Moreover, the reduced standard deviation of DICE scores under different missing modality settings indicates that our personalized approach consistently delivers superior segmentation results. This performance improvement stems from the enhanced generalization of our model, which is rooted in the learned ${\mathbb{X}_{h}}^{\prime}$.
5 Heterogeneous generalization: PET and CT
Given the differing imaging principles, the modality gap in heterogeneous generalization may be more pronounced than that in homogeneous generalization, making the former tasks more challenging. In this section, we evaluate our approach for heterogeneous generalization.
5.1 Pre-training for modality transfer
@@
Experimental settings. We utilize the AutoPET-II dataset [15] from the Automated Lesion Segmentation in PET/CT challenge for pre-training. The AutoPET-II dataset includes AC-PET and CT pairs, where the PET scans adopt FDG tracers, and their attenuation is corrected using the corresponding CT scans. Specifically, we divide the AutoPET-II dataset into training and testing sets. Similar to our approach for heterogeneous generalization, we adopt the Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) as evaluation metrics. In this section, we present the results of models that employ different combinations of the constraints and the set $\mathbb{O}$. We use both contrastive loss and the decomposition constraint as our baseline. Please refer to training details in Appendix D.
@@
Results. As shown in Tab. 4 and generated examples in Fig. 5, incorporating $\mathbb{O}$ with different combinations of constraints improves generation quality across most metrics. Specifically, using the constraints without $\mathbb{O}$ does not guarantee improvements, as discussed in Sec. 3.2.1. Ultimately, employing all constraints along with $\mathbb{O}$ yields the best average results across all translations, validating that our approach performs well in heterogeneous generalization settings. These results indicate that our method under the scope of personalization bridges the gap between structural and functional modalities. We validate the transferability of all these pre-train models in Sec. C.1, where additionally analysis are provided.
5.2 Tuning for segmentation

### v4

@@ 3.1.2 The connection between the constraints and global biological prior
It is important to note that the above constraints are closely interconnected, as they align with Eq. 1. After obtaining additional knowledge from $\mathbb{O}$, the invariance constraint ensures that the representations from each modality for a given individual are the same, such that ${\mathbb{X}^{i}_{h}}^{\prime}$ and ${\mathbb{X}_{h}}^{\prime}$ can be considered equivalent. Combined with the decomposition constraint, which enforces that ${\mathbb{X}_{h}}^{\prime}$ is shared for the generation of all possible modalities, ${\mathbb{X}_{h}}^{\prime}$ is thus able to generalize across modalities. Additionally, the equivariance and decomposition constraints implicitly maintain SO(3)-equivariance by satisfying the relation $\mathcal{D}\circ\mathcal{E}(\phi^{i}(X^{i}_{h});\theta)=\phi^{i}(\mathcal{D}\circ\mathcal{E}(X^{i}_{h};\theta))$ [^1] The SO(3) transformations are left-multiplication; they are expressed here in a simplified form, using $\phi^{i}(\cdot)$., where $\theta$ represents the model parameters after training. This ensures that geometric transformations are preserved in the latent features $z^{i}_{h}$, such that $\mathcal{F}(z^{i}_{h})=\phi^{i}$. The invariance constraint then requires that $z^{i}_{h}$ with geometric transformations can retrieve features from $\mathbb{O}$ to form ${\mathbb{X}^{i}_{h}}^{\prime}$, which remains invariant to any geometric transformation. This implicitly constrains $\mathbb{O}$ to contain comprehensive biological information, including other potential geometric transformations, thereby improving the $\mathbb{X}_{h}^{\prime}$ through $\mathbb{O}$ in approximating $\mathbb{X}_{h}$ and enhancing the robustness of ${\mathbb{X}_{h}}^{\prime}$.
3.2 Fine-tuning stage
@@
Empirically, we adopt the SwinUNETR architecture [19] as the backbone of the encoder $\mathcal{E}$, and implement the proposed components. The model is trained with $\mathcal{L}_{pre}$ during the pre-training phase; users have the option to either use the standard SwinUNETR by loading only our pre-trained encoder weights or to employ our proposed model structure with all pre-trained weights for downstream tasks. Notably, all modalities for a given individual, $X_{h}={X^{i}_{h},X^{j}_{h},...,X^{k}_{h}},i,j,...,k\in\mathcal{M}$, share the same encoder, with the encoder’s channel size set to match the number of modality types. The input volume size for all experiments is fixed at $96\times 96\times 96$. Further empirical details on how $\mathbb{X}_{h}$ is leveraged for homogeneous and heterogeneous generalization are provided in Secs. 4.1 and 4.2.
3.3 Theoretical analysis
Consider a mapping $g:\mathcal{H}\to\mathcal{M}$ that extracts modalities from the individual biological profile, where $g\in\mathcal{G}$ represents a set of candidate hypotheses. In modality generalization, our objective is to guarantee that the model effectively generalizes across diverse modalities $\mathcal{M}$ for previously unseen patients. To achieve this, the learning bound must account for the relationships between seen $\mathcal{H}^{\mathcal{S}}$ and unseen patients $\mathcal{H}^{\mathcal{U}}$. Inspired by the theoretical framework of domain adaptation (DA) risk bounds in [4], we derive the following bound for medical cross-modality generalization:
Theorem 3.2.
Let $R^{\mathcal{S}}$ and $R^{\mathcal{U}}$ denote the generalization errors on the modalities from the seen patient domain $\mathcal{D}^{\mathcal{S}}$ and the unseen patient domain $\mathcal{D}^{\mathcal{U}}$, respectively. For a given hypothesis $g\in\mathcal{G}$, the overall unseen risk is bounded by:
$\begin{split}R^{\mathcal{U}}(g)\leq R^{\mathcal{S}}(g)+d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{D}^{\mathcal{S}},\mathcal{D}^{\mathcal{U}})+\lambda,\end{split}$ (9)
where $d_{\mathcal{G}\Delta\mathcal{G}}$ is the $\mathcal{G}\Delta\mathcal{G}$-divergence between $\mathcal{D}^{\mathcal{S}}$ and $\mathcal{D}^{\mathcal{U}}$, and joint hypothesis $\lambda$ quantifies the inherent difficulty of aligning the seen and unseen domains (can be omitted).
The bound emphasizes the importance of minimizing the divergence between domains and achieving a low risk on the seen domain to ensure effective generalization to unseen domains. For a well-trained model on $R^{\mathcal{S}}(g)$ is minimized. the key focus shifts to $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{D}^{\mathcal{S}},\mathcal{D}^{\mathcal{U}})$. Considering modalities, $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{D}^{\mathcal{S}},\mathcal{D}^{\mathcal{U}})$ is defined as follows:
$\begin{split}d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{D}^{\mathcal{S}},\mathcal{D}^{\mathcal{U}})\!:=d_{\mathcal{G}\Delta\mathcal{G}}([\mathcal{H}^{\mathcal{S}}\oplus\mathcal{M}|\mathcal{H}^{\mathcal{S}}],[\mathcal{H}^{\mathcal{U}}\oplus\mathcal{M}|\mathcal{H}^{\mathcal{U}}]),\end{split}$
where $\oplus$ denotes that each domain combines the patient domain with its derived modalities. Upon rearranging, it follows that:
$d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{H}^{\mathcal{S}},\mathcal{H}^{\mathcal{U}})+d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}|\mathcal{H}^{\mathcal{S}},\mathcal{M}|\mathcal{H}^{\mathcal{U}}).$ (10)
Our approach minimizes $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}|\mathcal{H}^{\mathcal{S}},\mathcal{M}|\mathcal{H}^{\mathcal{U}})$ by initially learning $\mathbb{X}_{h}$ for each patient at the pre-training stage and then employing a fine-tuning stage to minimize $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{H}^{\mathcal{S}},\mathcal{H}^{\mathcal{U}})$. As shown in Fig. 2, while each modality uniquely visualizes a patient’s biological profile, the underlying generation principles remain unchanged. Thus, the mapping $g_{m}:\mathcal{H}\to\mathcal{X}_{m}$ that derives the modality-specific visualizations $m\in\mathcal{M}$ from the individual biological profile is consistent across individuals. Since direct access to a person’s anatomical structure is infeasible, we propose learning an invariant representation $\mathbb{X}_{h}$ for $h\in\mathcal{H}$ with a biological prior, allowing decomposition into various modalities. To ensure $\mathbb{X}_{h}$ effectively captures anatomical features, we enforce equivariance constraints to preserve geometric information. This learned representation augments the quality of $g_{m}$, improving its generalization ability by reducing $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}|\mathcal{H}^{\mathcal{S}},\mathcal{M}|\mathcal{H}^{\mathcal{U}})$. In the second fine-tuning stage, it is only necessary to mitigate $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{H}^{\mathcal{S}},\mathcal{H}^{\mathcal{U}})$, thus facilitating the transfer process.
Previous approaches ignore personal variations, opting instead to focus on minimizing the gap between modalities, as noted in $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}^{\mathcal{S}},\mathcal{M}^{\mathcal{U}})$. To facilitate a more effective comparison with our approach, we simplify this issue by positing that $d_{\mathcal{G}\Delta\mathcal{G}}$ is quantifiable via Kullback-Leibler (KL) divergence. Consequently, previous method aimed at minimizing $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}^{\mathcal{S}},\mathcal{M}^{\mathcal{U}})$ can be converted as follows:
$\begin{split}&d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}^{\mathcal{S}},\mathcal{M}^{\mathcal{U}}):=KL(\mathcal{M}^{\mathcal{S}}\|\mathcal{M}^{\mathcal{U}})\\ =&KL(\mathcal{H}^{\mathcal{S}}\|\mathcal{H}^{\mathcal{U}})+E_{\mathcal{H}^{\mathcal{S}},\mathcal{H}^{\mathcal{U}}}[KL(\mathcal{M}^{\mathcal{S}}|\mathcal{H}^{\mathcal{S}}\|\mathcal{M}^{\mathcal{U}}|\mathcal{H}^{\mathcal{U}})]\\ \geq&KL(\mathcal{M}^{\mathcal{S}}|\mathcal{H}^{\mathcal{S}}\|\mathcal{M}^{\mathcal{U}}|\mathcal{H}^{\mathcal{U}})=:d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{H}^{\mathcal{S}},\mathcal{H}^{\mathcal{U}})\;(\textbf{Ours}).\end{split}$
Our methodology narrows the gap identified in Eq. 10 compared to earlier methods, thereby demonstrating enhanced generalization and transferability capabilities.
4 Experiments
Overall experiment arrangement. Our experiments examine various generalization scenarios in two stages: generation pre-training and downstream fine-tuning. We assess pre-training by comparing generation quality to SOTA medical generation methods and evaluate fine-tuning by comparing with SOTA methods for each task, including benchmarking against SOTA medical self-supervised pre-training methods. More results and analysis are in Appendix G.
4.1 Homogeneous generalization: MRI
From scratch PSNR↑ NMSE↓ SSIM↑
SwinUNETR 29.1776 0.2196 0.8961
MS-SPADE 26.4225 0.0822 0.9086
Ours (PUIR) ($96\times 96\times 96$) 35.4378 0.0362 0.9587
Ours (PUIR) ($192\times 192\times 192$) 33.7418 0.0481 0.9502
Table 2: Modality transfer results of MRI on BRATS23: Comparison between the previous one-to-all modality transfer methods and ours for transfer between all four modalities. The averaged results of metrics for all validation samples across all modalities are listed. The best results are highlighted in blue. Please refer to Appendix Tab. 10 for more detailed results.
4.1.1 Pre-training stage: Modality transfer
Experimental settings. Following previous methods [27], we utilize the multi-modal brain tumor segmentation challenge 2023 (BRATS23) dataset [1, 35, 3, 2]. BRATS23 includes four structural MRI modalities (T1, T1ce, T2, and FLAIR) for each individual. Our model is tested on the BRATS23 validation set, which contains these four modalities for 219 individuals. We evaluate the quality of synthesis using peak signal-to-noise ratio (PSNR), normalized mean squared error (NMSE), and structural similarity index (SSIM) [53]. To provide comprehensive results, we separately compare the translation results for T1 $\to$ T2 and T2 $\to$ FLAIR, as some previous methods are only capable of single-modality transfer. These include both 2D and 3D generation methods, as shown in Tab. 1. Additionally, we employed SwinUNTER for multi-modality translation comparisons. All evaluations were performed on 3D volumes; for the 2D methods, synthesized target images were stacked to form a 3D volume for comparison. Please refer to model training details to Appendix F.
@@ 4.1.1 Pre-training stage: Modality transfer
SwinUNETR pre-training 86.55 3.49 89.22 88.18 1.38 86.69 3.15 84.04 4.91
Ours (PUIR) 87.63 3.25 91.19 89.49 1.82 87.45 1.03 85.17 3.93
Table 3: Missing modality segmentation results of MRI on BRATS18: Comparisons between previous SOTA methods and ours. MN: Number of how many missing modalities; FM: Full Modality. We report mean and standard deviations of DICE results under the MN. The best DICE results are highlighted in blue. Please refer to Appendix Tab. 11 for more details.
@@
Figure 4: Generated samples on AUTOPET-II.
@@
Results. Tab. 1 exhibits the transfer results in comparison between our and previous one-to-one modality transfer methods. Following previous of T1 $\to$ T2 and T2 $\to$ Flair. Our approach significantly surpasses previous 2D and 3D generation methods, including single- and multi-modality translation methods. Specifically, our approach exceeds current SOTA diffusion-based methods, such as 2D-based MD-Diff and 3D-based MS-SPADE. In terms of one-to-all multi-modality translation, as Tab. 2 and additional results in the Appendix show, our approach with various volume sizes ($96$ and $192$) performs better than MS-SPADE and SwinUNETR across all metrics under all settings. Moreover, our method significantly improves the SSIM, indicating a better anatomy structure obtained by our approach. These results indicate that the $\mathbb{X}_{h}$ Hypothesis is plausible for homogeneous generalization, and our personalized approach is able to obtain its approximation. Please refer to more analysis on $\mathbb{O}$ in Sec. G.1.
Figure 5: T-SNE visualizations on embeddings of SwinUnetr and ours trained on BRATS23 for one-to-all modalities generation.
More analysis on personalization. Fig. 5 also shows that combining the prior and constraints results in more scattered embeddings at the individual level compared to the method without them. Tab. 6, Tab. 3, and the quantitative results in the paper further confirm the advantages of personalization.
4.1.2 Fine-tuning stage: Missing modality segmentation
@@
Experimental settings. To validate the generalization ability of the pre-trained model, we fine-tune the model obtained from Sec. 4.1.1 on the BRATS18 [35] from the Multimodal Brain Tumor Segmentation Challenge. Similar to BRATS23, BRATS18 also consists of the same four structural modalities. We employ the Dice similarity coefficient (DICE) as the metric for evaluation. For a fair comparison, we follow data splits of [42] and reproduce the results of previous methods [14, 57, 46, 31, 42] on these splits by using their released code and following their original settings  [^2] Though we tried our best, it can be noticed some reproduced results are lowered than their reported results in their original paper. It should be clarified that our results also exceed those reported results. However, for a comprehensive study, we mainly report our reproduced results. . We also reproduce the results by using the SwinUNETR pre-training method [44], employing their official code to train the model on the same datasets with identical training epochs, learning rates, and other hyperparameters as ours. See additional experimental details in Appendix F.
@@
Results. Tab. 3 presents the segmentation results of our approach compared to previous methods. We also compute the standard deviation of DICE scores under various missing modality settings, which highlights the robustness of models. Notably, our approach outperforms previous methods in most missing modality scenarios, particularly when the number of missing modalities is large (e.g., NM=3). Meanwhile, our approach also obtain a relatively low standard deviation for most settings, showing its robustness to various modalities missing scenario. Compared with SwinUNETR pre-training, our methods yield consistently better DICE with comparable standard deviation. This performance improvement stems from the enhanced generalization of our model, which is rooted in the learned ${\mathbb{X}_{h}}^{\prime}$.
4.2 Heterogeneous generalization: PET and CT
4.2.1 Pre-training stage: Modality transfer
@@
Experimental settings. We utilize the AutoPET-II dataset from the Automated Lesion Segmentation in PET/CT challenge [17] for pre-training. The AutoPET-II dataset includes AC-PET and CT pairs, where the PET scans adopt FDG tracers, and their attenuation is corrected using the corresponding CT scans. Specifically, we divide the AutoPET-II dataset into training and testing sets. Similar to our approach for heterogeneous generalization, we adopt the Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) as evaluation metrics. In this section, we present the results of models that employ different combinations of the constraints and the set $\mathbb{O}$. We use both contrastive loss and the decomposition constraint as our baseline. Please refer to training details in Appendix F.
@@
Results. As shown in Tab. 4 and generated examples in Fig. 4, incorporating $\mathbb{O}$ with different combinations of constraints improves generation quality across most metrics. Specifically, using the constraints without $\mathbb{O}$ does not guarantee improvements, as discussed in Sec. 3.1.2. Ultimately, employing all constraints along with $\mathbb{O}$ yields the best average results across all translations, validating that our approach performs well in heterogeneous generalization settings. These results indicate that our method under the scope of personalization bridges the gap between structural and functional modalities. We validate the transferability of all these pre-trained models in Appendix E, where additional analyses are provided.
4.2.2 Fine-tuning stage: Segmentation